OpenStudy (anonymous):

how do i find the derivative of y= 1/sin^-1(2x)

OpenStudy (anonymous):

what is your quesiton about the problem?

OpenStudy (abb0t):

\(\sf \color{}{\frac{1}{arcsin(2x)}}\) is that it?

OpenStudy (anonymous):

i know i have to use the quotient rule but i dont understand what (sin^-1)^2 means. i dont know how to simplify after plugging in according to the quotient rule

OpenStudy (abb0t):

Wait, is it \(\sf \color{}{\frac{1}{[arcsine(x)]^2}}\) or what? can u rewrite it usng latex?

OpenStudy (shamil98):

Use the quotient rule. Remember: (arcsinx)' = 1/√(1-x^2)